OpenStudy (anonymous):

solve for x and y 3/x=2/5 y/4=y-1/3 x= y=

OpenStudy (anonymous):

3/x=2/5 Cross-multiply them, and 15/2=x y/4=y-1/3 1/3=y-y/4 1/3=3y/4 4=9y 4/9=y

jhonyy9 (jhonyy9):

15=2x --- x= 15/2 3y=4y-4 4=y

OpenStudy (anonymous):

thank u guyz :)

OpenStudy (anonymous):

No problem =]

OpenStudy (anonymous):

:)